OpenStudy (anonymous):

whats the x intercept and the y intercept of x + y = -3

OpenStudy (anonymous):

Make y=0 and find x

OpenStudy (anonymous):

x+0 =-3

OpenStudy (anonymous):

The x-intercept is (-3,0)

OpenStudy (tkhunny):

Ever meet the Intercept Form? Just divide by -3, in this case. \(\dfrac{x}{-3} + \dfrac{y}{-3} = 1\) Both intercepts are well-identified in this form.